OpenStudy (anonymous):

Find the slope of a line parallel to 5x + 2y = 6. A. two fifths B. negative two fifths C. negative five halves D. five halves

OpenStudy (alexandervonhumboldt2):

parallel lines have same slope in form Ax+By=c the -a/b is slope

OpenStudy (calculusxy):

this is in standard form: \(\large ax + by = c\) you need to put it into slope-intercept form \(\large y = mx + b\) \(\large \color{red}{5x - 5x + 2y = 6 - 5x}\) \(\large \color{green}{2y = 6 - 5x}\) \(\large \color{green}{ y = 3 - \frac{5}{2}x}\) \(\rightarrow\) \(\large \color{green}{y = -\frac{5}{2}x + 3}\) parallel lines have the same slope so look for the \(\large m\) in this equation